Translate the following statement into an algebraic equation. let x represent the number. 1 less than 15 times a number is 9 times that same number.
Thank you for any help I may be given.
D Jack
sgtdjack@yahoo.com

Found 2 solutions by macston, fractalier:
Answer by macston(5194)

.
15x-1=9x
.
6x-1=0
6x=1
x=1/6
.

Answer by fractalier(6550)

Okay, sarge, thanks for your service (I was Army 1972-74)...
15x - 1 = 9x
Remember that the word "is" means "equals."
Should you wish to solve this, we could subtract 15x from both sides and get
-1 = -6x
then divide by -6 and get
x = 1/6